OpenStudy (anonymous):

express as a sum or difference of logarithms without exponents. *problem in post below*

OpenStudy (anonymous):

\[\log_{c5}\sqrt{\frac{ x ^{2} }{ y ^{5}z ^{7}}} \]

OpenStudy (anonymous):

answer asks "what is the equivalent sum or difference of logarithms?"

OpenStudy (anonymous):

can't you try this: log a/b = loga - logb ?

OpenStudy (anonymous):

the thing i'm confused about is the logc5. im not good with log. never have been and have no idea how to do anything with it. ive tried guides online but everything just made me more confused

hartnn (hartnn):

actually c5 doesn't make sense to me either. ignoring that we can write \(\huge{\sqrt{\frac{x^2}{y^5z^7}}=\frac{x}{y^{5/2}z^{7/2}}}\) clear with this first step ??

OpenStudy (anonymous):

yeah. just taking the square off the first one and dividing the bottom squares by two right?

OpenStudy (anonymous):

so we can take the square root off

hartnn (hartnn):

more specifically using this : \(\sqrt{a^n}=a^{n/2}\) ok, now log properties: 1st : \(log (A/B)=log A -log B \) so that would make \(log x - log (y^{5/2}x^{7/2})\) ok with 2nd step ??

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

so would that be my answer?

hartnn (hartnn):

another log property : logAB = log A+log B so \(log(y^{5/2}z^{7/2})=log(y^{5/2})+log(z^{7/2}) \) this is 3rd step and 4th step is remaining.

OpenStudy (anonymous):

so would i replace \[\log x - \log(y ^{5/2}x ^{7/2}) \]

OpenStudy (anonymous):

with \[logx - \log(y ^{5/2}) + \log(y ^{7/2)}\]

OpenStudy (anonymous):

\[\log(z ^{7/2}) i mean\]

hartnn (hartnn):

its \(logx - \log(y ^{5/2}) - \log(y ^{7/2)}\)

hartnn (hartnn):

now last step, use log (a^n) = n log a

OpenStudy (anonymous):

ah i see where i messed up there

OpenStudy (anonymous):

where do i get the n from?

hartnn (hartnn):

log (y^(5/2)) = (5/2) log y ok ? can u do the same with z ?

OpenStudy (anonymous):

(7/2) log z

hartnn (hartnn):

yup,so whats final answer?

OpenStudy (anonymous):

logx - (5/2) log y - (7/2) log z?

hartnn (hartnn):

\(log x - (5/2) log y -(7/2) log z.\) good work :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome :)